OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9.

OpenStudy (anonymous):

Hey, you are here, we are searching you at your previous post. Are you playing hide and seek with us? :P

OpenStudy (anonymous):

Okay slope of equation of line in standard form ie \(ax + by = c\) is given by: \[\large \color{green}{Slope = \frac{- (x \; \; coefficient)}{(y \; \; coefficient)}}\]

OpenStudy (anonymous):

Can you tell what is the \(x\) coefficient here?

OpenStudy (anonymous):

\[so -4/5  and 9/5 ?\]

OpenStudy (anonymous):

-4/5 is good for slope.. Recheck for your y intercept part..

OpenStudy (anonymous):

-4/5 and -/9/5?

OpenStudy (anonymous):

For finding y-intercept: 1. Take 4x to other side by subtracting 4x from both the sides. 2. Divide by 5 to get equation in the form of y = mx + c, where c= y-intercept..

OpenStudy (anonymous):

yep, -4/5 and -9/5 are good.. :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

\(\dagger\)

OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line?

OpenStudy (anonymous):

You know equation of line in Slope intercept form is given as: \(y = mx + b\), m is slope and c is y-intercept..

OpenStudy (anonymous):

If a line cuts y-axis at a point, then the y-coordinate of that points is the y-intercept and you are given slope, m = 45 already.. Can you write the equation now?

OpenStudy (anonymous):

What is the y-coordinate of the point (0,1) ??

OpenStudy (anonymous):

5y - 4x = 5?

OpenStudy (anonymous):

Nope..